# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v1

@@ Abstract.
For connected reductive groups, we prove that the cohomology of the structure sheaf and the canonical sheaf for compactified Deligne–Lusztig varieties associated to a tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements depends only on the $F$-support of $(w_{1},\ldots,w_{r})$.
@@ 1.1. Background
Let $q$ be a power of a prime $p$ and let $k$ be an algebraic closure of the finite field $\mathbb{F}_{q}$. Let $G_{0}$ be a connected reductive $\mathbb{F}_{q}$-group and $G$ its base change to $k$. Denote by $F:G\to G$ the corresponding $q$-Frobenius. The group of fixed points $G^{F}:=G(\mathbb{F}_{q})$ of $F$ is also known as a finite group of Lie type in the literature.
Let $T_{0}\subset B_{0}$ be a maximal torus and a Borel subgroup of $G_{0}$ and $T\subset B$ their base change to $k$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$, and there is a smooth compactification $\overline{X}(w)$ associated to each reduced expression of $w$. In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic cohomology. However, when $G={\rm GL}_{2}$ and $w$ is the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
The simplicity in the situation of $\overline{X}(w)$ for $G={\rm GL}_{2}$ does not carry over to higher dimensions. The smooth compactification $\overline{X}(w)$ is constructed in loc. cit. as the transversal intersection between the graph of the Frobenius and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen. Unlike the desingularisation of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration. Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
In [Wan], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of $G={\rm GL}_{n}$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S|s\leq w\}$. As a consequence of our main theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lustig variety $X(w)$ for ${\rm GL}_{n}$.
@@ 1.2. Main results
The aim of this paper is to extend the main theorem of [Wan] to all connected reductive groups $G_{0}$.
@@
In this general setting, we encounter a special class of elements in $W$ which enjoy nice properties [GKP], [HN]. Namely, they are the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2). In any given $F$-conjugacy class $C$, we denote by $C_{\rm min}$ the subset of minimal length elements. For example, if $G$ is ${\rm GL}_{n}$ or an inner form of ${\rm GL}_{n}$, such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
@@
We show that the study of the cohomology of the structure sheaf and canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of some minimal length element in an $F$-conjugacy class.
Theorem 1.1.
@@
Let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(w)},\mathcal{O}_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\mathcal{O}_{{\overline{X}(x)}}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}(w)},\omega_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\omega_{{\overline{X}(x)}}\Big{)}$
for all $k\geq 0$. Moreover, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(x)},\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
for all $k\geq 0$, where $I$ is the $F$-support of $x$, $P_{I}$ and $L_{I}$ are the associated standard parabolic and Levi subgroup, and $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
@@
One should note that the $F$-conjugacy class $C$ in the above theorem may not contain $w$. However, for any given $w$, one can determine explicitly the minimal length elements in this $F$-conjugacy class by following our recipe given in the proof of 5.1.
@@
Smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements cf. 2.2. These notions also make the construction of partial compactifications of Deligne–Lusztig varieties more straightforward and uniform. One can always recover statements for classical Deligne–Lusztig varieties by setting $r=1$.
@@
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BR]. We also show that they have $F$-rational singularities and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of partial compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
@@
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we prove that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{1},\ldots,w_{r}$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and canonical sheaf on the associated compactified Deligne–Lusztig varieties.
@@
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In 4.6, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties. Furthermore, we deduce that the cohomology of the structure sheaf (resp. canonical sheaf) for $\overline{X}(w)$ and $\overline{X}(F_{1}^{m}(w))$ are isomorphic for all $m\geq 0$, where $F_{1}$ is a partial Frobenius morphism (2.5). The key observation in this section is that the partial Frobenius morphism $F_{1}$ induces an automorphism of the Weyl group, even though it has no well-defined inverse as a morphism of schemes.
In Section 5, we begin with proving the statement in 5.1 under the assumption that $F(S)=S$. The remaining cases come from Suzuki and Ree groups, which are treated in Section 5.2. Finally, the theorem is proved in full generality in Corollary 5.3.
Acknowledgements
@@ 2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. By adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and smooth, respectively.
@@ 2.1. Definitions and basic properties
Let $G$, $B$, $T$, $F:G\to G$, $W$, $S$ be as in Section 1.1. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
@@
We use such tuple of elements $w_{1},\ldots,w_{r}\in W$ for the definition of partial compactifications. They will in particular be useful for avoiding ambiguity in having different reduced expressions of a Weyl group element.
Definition 2.1.
Let $w_{1},\ldots,w_{r}\in W$. We define the compactified orbit associated to $(w_{1},\ldots,w_{r})$ as the $k$-scheme
$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})},$
where for each $i=1,\ldots,r$, $\overline{O(w_{i})}$ is the Zariski closure of $O(w_{i})$ in $G/B\times G/B$.
The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ can be viewed as an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$. Thus $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is reduced, separated, normal, Cohen Macaulay $k$-schemes locally of finite type. The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G$-equivariant embedding $O(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}})$ cf.[DMR, Proposition 2.2.16]. Moreover, there is a stratification
$O(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}O(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
Denote by $F:G/B\to G/B$ the map induced by the Frobenius $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
Definition 2.2.
@@
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. By Lemma 2.3 below, $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is smooth. To simplify the notation, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
@@
Lemma 2.3.
Let $w_{1},\ldots,w_{r}\in W$. The Deligne–Lusztig variety $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
Proof.
In the proof of [DL, Lemma 9.11], Deligne and Lusztig showed that the graph $\Gamma_{F}$ of the Frobenius is transversal to any smooth $G$-equivariant scheme $Y\to G/B\times G/B$. Here let $Y=O(w_{1},\ldots,w_{r})$ and $Y=O(\underline{w_{1}},\ldots,\underline{w_{r}})$, which are $G$-equivariant schemes over $G/B\times G/B$. By construction, $O(w_{1},\ldots,w_{r})$ is an iterated fibration over $G/B$ with fibres being Schubert cells, which are affine spaces. The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibration over $G/B$ with fibres being Schubert varieties $\overline{{B}w_{i}{B}}/{B}$. Hence the smoothness of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ follows. ∎
Remark 2.4.
In general, compactified Deligne–Lusztig varieties are $F$-rational and thus pseudo-rational, normal, and Cohen–Macaulay. We will give a proof for this in 2.13.

### v2

@@ Abstract.
For connected reductive groups together with a Frobenius root $F$ that preserves the set of simple reflections, we show that for every tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements there exists a minimal length element in an $F$-conjugacy class in $W$ such that the cohomology of the structure sheaf and the canonical sheaf for the associated compactified Deligne–Lusztig varieties are isomorphic respectively.
@@ 1.1. Background
Let $q$ be a power of a prime $p$ and let $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$. Let $k$ be an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$. Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$. Let $G$ be the base change of $G_{0}$ to $k$ and let $F:G\to G$ be an endomorphism such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$. The group of fixed points $G^{F}$ is known as a finite group of Lie type.
Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$. Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$.
In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero. The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart. However, for $G^{F}={\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
The simplicity in the situation of $\overline{X}(w)$ for ${\rm GL}_{2}(\mathbb{F}_{q})$ does not carry over to higher dimensions. The smooth compactification $\overline{X}(w)$ is constructed in loc. cit. as the transversal intersection between the graph of $F$ and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen. Unlike these desingularisations of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration. Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
In [Wang22], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$. As a consequence of this theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lusztig variety $X(w)$ for ${\rm GL}_{n}(\mathbb{F}_{q})$.
@@ 1.2. Main results
The aim of this paper is to extend the main theorem of [Wang22] to all pairs $(G,F)$ for which $F(S)=S$. The assumption on $F$, for example, is satisfied when $G$ is split over $k_{0}$, or when $G^{F}$ is the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
If $G^{F}$ is a finite simple group of Lie type arising from a pair $(G,F)$ as in in Section 1.1, then the classification of such groups implies that $F(S)=S$ is satisfied. In fact, $F$ would be the Frobenius endomorphism (i.e. $d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section 5.2).
@@
In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2), which is a special class of elements in $W$ with nice properties [GeckKimPfeifferMinlength, Theorem 2.6], [HeNieminlength, Theorem 3.1]. For example, if $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$ or the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$, then such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
@@
We show that the study of the cohomology of the structure sheaf and the canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of a minimal length element in an $F$-conjugacy class in $W$.
Theorem 5.1.1.
@@
Assume $F(S)=S$ and let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left({\overline{X}(w)},\mathcal{F}_{{\overline{X}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}(x_{0})},\mathcal{F}_{{\overline{X}(x_{0})}}\right),$
where $\mathcal{F}\in\{\mathcal{O},\omega\}$ is either the structure sheaf or the canonical sheaf.
@@
Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$. However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem 5.1.1. We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
@@
On the other hand, smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf. Definition 2.1.1, [DMR, Definition 2.3.2]. This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform. For example, setting $r=1$ yields the Zariski closure $X(\underline{w_{1}})=\overline{X(w_{1})}$ of $X(w_{1})$ in $G/B$. Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
@@
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BonnafeRouquierOntheirred]. We also show that they have $F_{p}$-rational singularities (Section 2.4) and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
@@
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
@@
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In Proposition 4.1.4, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties.
Finally, we prove Theorem 5.1.1 and extend the result to arbitrary compactified Deligne–Lusztig varieties of the form $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in Section 5.
@@ 2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. This construction has been treated in [DL, §9] and [DMR, §2]. We review here the definitions and properties that are relevant for our context. Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and show that they are $F_{p}$-rational.
@@ 2.1. Definitions and basic properties
Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section 1.1. In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
@@
Moreover, we denote the fibre product of the Zariski closures $\overline{O(w_{i})}\subseteq G/B\times G/B$ of the orbits by
$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})}.$
We refer to them as the compactified orbit associated to $(w_{1},\ldots,w_{r})$ in the rest of the article. The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf. [Jantzen, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type. A proof of normality can be found in [DMR, Lemma 2.2.13]. The Cohen–Macaulay property behaves well under base change [stacks-project, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [Hashimoto]).
Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph. Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([DigneMichelBook, Proposition 8.1.2]), separated ([GoerWedhAlgGeom1, Proposition 9.13]), and purely inseparable ([EGA1, (3.7.6)]). Moreover, the induced map $F:W\to W$ is a group automorphism [DigneMichelBook, Lemma 4.2.21].
Definition 2.1.1.
@@
For any $w_{1},\ldots,w_{r}\in W$, the $k$-scheme $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf. [DMR, Proposition 2.3.5, 2.3.6 (i)]. In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
@@
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf. [DMR, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth. To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.